DROP DOWN OPTIONS

1. (28%,19.50%,25%)

2. (19.50%, 20.80%,29.20)

3. (25%, 15%,16.67%)

4. (16.67%, 29.20%, 18.15%

2. Calculating tax liabilities Calculating Taxable Income Taxable income Rate $1 $9,075 10.0% $9,076-$36,900 15.0% 25.0% $36,What is Chloes tax liability? What is Chloes top marginal tax rate? What is Chloes average tax rate? Clarks Tax Scenario

2. Calculating tax liabilities Calculating Taxable Income Taxable income Rate $1 $9,075 10.0% $9,076-$36,900 15.0% 25.0% $36,901-$89,350 $89,351 _ $186,350 28.0% $186,351-$405,100 33.0% $405,101-$406,750 35.0% 39.6% Over $406,751 Using the previous tax table, compute the tax liability for each individual in the scenarios presented, rounding the liability to the nearest dollar. In addition, use the dropdown lists to identify the marginal tax rate and average tax rate for both Chloe and Clark. Chloe's Tax Scenario Chloe is a young professional who lives in downtown San Francisco. She has taxable income of $94,735 as an advertising account executive for Charles Underwood Agency Inc.
What is Chloe's tax liability? What is Chloe's top marginal tax rate? What is Chloe's average tax rate? Clark's Tax Scenario Clark is a young professional who lives in downtown Seattle. He has taxable income of $60,500 as a medical sales rep for Frasier-dool Inc. What is Clark's tax liability?s What is Clark's top marginal tax rate? What is Clark's average tax rate?
Answer #1

Answer:;

Given

Choles Taxbale income =$94735

Since given income falls between 89351-186350 so marginal tax rate =28%

Tax Liability = 9075*10%+(36900-9075)*15%+(89350-36900)*25%+(94735-89350)*28%=$19701.55

Marginal tax rate =28%

Average tax rate = tax liability/ Taxable income = 19701.55/94735=20.80%

Clarks Taxable income =$60500

Since given income falls between 36901-89350 so marginal tax rate =25%

Tax Liability = 9075*10%+(36900-9075)*15%+(60500-36900)*25%=$10981.25

Marginal tax rate =25%

Average tax rate = tax liability/ Taxable income = 10981.25/60500=18.15%